Good Morning,
PLEASE Help learn how to solve the below methods.
1.Solve using the elimination methods
x + 6y = 44
-x + 4y = 16

Thanking you in advance Betty

Found 2 solutions by Alan3354, fractalier:
Answer by Alan3354(69443)

x + 6y = 44
-x + 4y = 16
----------------- Add
10y = 60
etc

Answer by fractalier(6550)

Hi Betty...we can add these equations just as they are and we get
x + 6y = 44
-x + 4y = 16
--------------
10y = 60
Now divide by 10
y = 6
Now substitute into the first equation to find x...
x + 6(6) = 44
x + 36 = 44
x = 8